OpenStudy (anonymous):

Solve: 0 = 4x2 − 8x − 5.

OpenStudy (anonymous):

4x²-8x-5 = 0 x²-2x-5/4 = 0 (x -1)² - 1 - 5/4 = 0 {completing the square} (x -1)² - 9/4 = 0 (x -1)² = 9/4 x -1 = ±√(9/4) x = 1 ±√(9/4) x = 1±(3/2) x = 1+(3/2) or 1-(3/2) x = 5/2 or -1/2

OpenStudy (anonymous):

@simplybeyoutiful got any more question

OpenStudy (anonymous):

i just dont really get how to do it

OpenStudy (anonymous):

there are many ways of going this

OpenStudy (anonymous):

i think this is the easiest way

OpenStudy (cwrw238):

oooppps i'm wrong it will factor if you put 4 = 2*2 and -5 = -5 * +1 and combine these factors (one from each) to get -8 we can factor 2 * -5 + 2 * 1 = -8 giving (2x - 5)(2x + 1) = 0 so 2x - 5 = 0 x = 5/2 or 2x + 1 = 0 x = -1/2 solution st is 5/2, -1/2

OpenStudy (anonymous):

okay thank you, but @laxplayer how did you do it?

OpenStudy (anonymous):

Okay i will show and ur welcome @simplybeyoutiful

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Assume that your function can be written in the form (ax+b)(cx+d) = acx² +adx + bcx + bd = (ac)x² +(ad+ bc)x + bd Compairing this to 4x²-8x-5 = 0 shows that we want to find a,b,c,d such that Ac = 4 Ad + bc = -8 Bd = -5 Letas assume a,b,c,d are all integers and that a > 0 From ac = 4 we have the following options A = 1 and c = 4 A = 2 and c = 2 (we can ignore a = 4 and c = 1 as we could just rename the variable) From Bd = -5 we have the following options B = 1 and d = -5 B = -1 and d = 5 B = 5 and d = -1 B = -5 and d = 1 So this gives us the following options (x + 1)(4x - 5) (x - 1)(4x + 5) (x + 5)(4x - 1) (x - 5)(4x + 5) (2x + 1)(2x - 5) (2x - 1)(2x + 5) (2x + 5)(2x - 1) ******this is the same as two above so can be ignored (2x - 5)(2x + 1) ******this is the same as two above so can be ignored Now we can just expand each of these to see which one if right! (x + 1)(4x - 5) = 4x² - 5x + 4x - 5 = 4x² - x - 5 (x - 1)(4x + 5) = 4x² + 5x - 4x - 5 = 4x² + x - 5 (x + 5)(4x - 1) = 4x² - x + 20x - 5 = 4x² + 19x - 5 (x - 5)(4x + 5) = 4x² + x - 20x - 5 = 4x² - 19x - 5 (2x + 1)(2x - 5) = 4x² - 10x + 2x - 5 = 4x² - 8x - 5 (2x - 1)(2x + 5) = 4x² + 10x - 2x - 5 = 4x² + 8x - 5 You should have spotted that the second to last one was right! so 4x²-8x-5 = 0 Becomes (2x + 1)(2x - 5) = 0 But for xy = 0 either x=0 or y=0 (or both) So either (2x + 1) = 0 → 2x = -1 → x = -1/2 Or (2x -5) = 0 → 2x = 5 → x = 5/2

OpenStudy (anonymous):

i think i might get it haha

OpenStudy (anonymous):

@simplybeyoutiful hope this helps